Sioux Corporation is estimating the following sales for the first four months of next year:

January: $260,000

Febraury: $230,000

March: $270,000

April: $320,000

Sales are normally collected 60% in the month of slae, 35% in the month following the sale, and the remaining 5% being uncollectible. Based on this information, how much cash should Sioux expect to collect during the month of April?

A.) $286,500

B.) $320,000

C.) $192,000

D.) $94,500
Answer #1
Concepts and reason

Cash collection schedule: This schedule is prepared for the estimation of the cash collection for the period. This includes probable cash from the sources. This is a very important schedule. This schedule determines the expected cash inflow from the operating activities.

Fundamentals

Cash collection from different periods is calculated by multiplying the credit sales value with expected cash collection percentage. All expected cash calculations for the period are added together to find the total expected cash collection for the period.

Credit sale for March is $200,000. Cash collected from the March month credit sales, during April is 35% of total credit sales.

Determine the expected cash collection from March sale as follows: -

CashcollectionfromMarchsale=Marchsale×35%=$270,000×35%=$94,500\begin{array}{c}\\Cash\,\,collection\,\,from\,\,March\,\,sale = March\,\,sale\,\, \times 35\% \\\\ = \$ 270,000\, \times 35\% \\\\ = \$ 94,500\\\end{array}

Credit sale for April is $320,000. The expected cash collection for April sale is 60% of the total sales. The expected cash collection from the March sale is $94,500.

Determine the cash collection during April as follows: -

Cashcollected=CashcollectedforMarch+CashcollectedforAprilSale=$94,500+($320,000×60%)=$286,500\begin{array}{c}\\Cash\,\,collected = Cash\,\,collected\,\,for\,March + Cash\,\,collected\,\,for\,\,April\,\,Sale\\\\ = \,\$ 94,500 + \,(\$ 320,000\,\, \times \,\,60\% )\\\\ = \$ 286,500\\\end{array}

Therefore, the total cash collected during April is $286,500.

Ans:

The total cash collected during April is $286,500.